### v2

# SBoRA: Low-Rank Adaptation with Regional Weight Updates

Abstract
This paper introduces Standard Basis LoRA (SBoRA), a novel parameter-efficient fine-tuning approach for Large Language Models that builds upon the pioneering works of Low-Rank Adaptation (LoRA) and Orthogonal Adaptation. SBoRA further reduces the computational and memory requirements of LoRA while enhancing learning performance. By leveraging orthogonal standard basis vectors to initialize one of the low-rank matrices, either $\mathbf{A}$ or $\mathbf{B}$, SBoRA enables regional weight updates and memory-efficient fine-tuning. This approach gives rise to two variants, SBoRA-FA and SBoRA-FB, where only one of the matrices is updated, resulting in a sparse update matrix $\mathrm{\Delta}\mathbf{W}$ with a majority of zero rows or columns. Consequently, the majority of the fine-tuned model’s weights $(\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain unchanged from the pre-trained weights. This characteristic of SBoRA, wherein regional weight updates occur, is reminiscent of the modular organization of the human brain, which efficiently adapts to new tasks. Our empirical results demonstrate the superiority of SBoRA-FA over LoRA in various fine-tuning tasks, including commonsense reasoning and arithmetic reasoning. Furthermore, we evaluate the effectiveness of QSBoRA on quantized LLaMA models of varying scales, highlighting its potential for efficient adaptation to new tasks. Code is available at https://github.com/cityuhkai/SBoRA
Keywords:
Large Language Models Parameter-Efficient Fine-Tuning LoRA
1 Introduction
Large language models (LLMs) and diffusion models have become essential components of natural language processing [1, 25] and multimodal AI applications [14, 16]. Full fine-tuning (FFT) has been shown to significantly improve their performance in various downstream tasks [28] or introduce new concepts. However, fine-tuning pre-trained models, especially LLMs, by updating all parameters is computationally costly. As depicted in Fig. 1(a), FFT involves directly updating the high-dimensional pre-trained weight matrix $\mathbf{W}_{0}$.
To address this issue, parameter-efficient fine-tuning (PEFT) methods [9] have gained momentum, focusing on updating only a small fraction of parameters, like adapter [7, 9, 19], prompt tuning [13, 21, 26]. Among these methods, Low-Rank Adaptation (LoRA) has emerged as a groundbreaking technique for LLMs. LoRA [10] introduces a parallel low-rank adapter to the weights of linear layers as shown in Fig. 1(b), reducing memory overhead and computational costs during fine-tuning. However, there are still limitations to LoRA, including activation memory consumption and a performance gap compared to FFT. These limitations have led to the development of many LoRA variants [2, 3, 4, 6, 17, 22, 23, 27, 29, 31, 32].
Figure 1: Four fine-tuning strategies: (a) Full Fine-Tuning (FFT), (b) LoRA, (c) SBoRA-FA, and (d) SBoRA-FB.
In this work, we introduce Standard Basis LoRA (SBoRA), a novel approach that further reduces the computational and memory requirements while enhancing the learning capacity. SBoRA achieves this by selectively updating specific rows or columns of the pre-trained weight matrix $\mathbf{W}_{0}$, preserving most of the base model’s weights. The resulting update matrix $\mathrm{\Delta}\mathbf{W}=\mathbf{BA}$ mainly consists of zero rows or columns, indicating that most of the fine-tuned model’s weights $(\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain identical to the pre-trained weights. This localized learning process parallels the modular organization of the human brain, where specific cognitive functions are localized in distinct regions [5]. For example, the hippocampus is responsible for episodic memory [30] while the prefrontal cortex is crucial for working memory and decision-making [12]. By mirroring the modular functionality of the human brain, SBoRA is expected to improve knowledge retention and facilitate efficient adaptation.
Let’s assume that the pre-trained weight matrix $\mathbf{W}_{0}$, the projection-down matrix $\mathbf{A}$, and the projection-up matrix $\mathbf{B}$ are given. SBoRA adopts a unique approach, utilizing orthogonal standard basis vectors (one-hot vectors) to construct its projection matrices. These matrices, designated as $\mathbf{A}_{sb}$ for SBoRA-FA and $\mathbf{B}_{sb}$ for SBoRA-FB, are illustrated in Figures 1(c) and 1(d), respectively. This efficiency stems from the elimination of the need to store coefficients for one of the projection matrices, resulting in approximately 50% reduction in memory usage. In SBoRA-FA, besides the frozen weight matrix $\mathbf{W}_{0}$, the $\mathbf{A}_{sb}$ is fixed and only updates $\mathbf{B}$ during fine-tuning. The weight changes ($\mathrm{\Delta}\mathbf{W}=\mathbf{BA}_{sb}$) reside in a low-rank subspace spanned by $\mathbf{A}_{sb}$, and the projection-down transformation can be efficiently achieved by selecting a subset of input vector samples. SBoRA-FB operates similarly, using a pre-defined standard basis projection-up matrix $\mathbf{B}_{sb}$ and updating only matrix $\mathbf{A}$. Both SBoRA-FA/FB align with LoRA, reducing computational overhead during fine-tuning and introducing no extra inference latency.
We extensively validated SBoRA across diverse tasks, including evaluations on floating-point 16-bit precision models and quantized models in the QLoRA framework [4]. For 16-bit models, we focused on commonsense reasoning and arithmetic reasoning tasks. Quantized models were evaluated using the MMLU benchmarks [8]. Experimental results consistently demonstrate SBoRA’s superiority over other PEFT baselines. For instance, with similar number of trainable parameters (SBoRA: rank 64, LoRA/DoRA: rank 32), SBoRA-FA demonstrates significant improvements in commonsense reasoning (+2.9%/1.7% on LLaMA-7B/LLaMA3-8B) compared to LoRA, and arithmetic reasoning (+2.8%/+2.0% on LLaMA-7B/LLaMA3-8B) compared to DoRA. Furthermore, with approximately half the trainable parameters (QSBoRA/QLoRA/QDoRA: rank 64), QSBoRA-FA exhitbs notable enhancements on MMLU benchmarks, such as +4.3%/+6.9% on quantized LLaMA-13B/LLaMA3-8B compared to QLoRA.
The summary of our contributions is as follows:
  • Introduces regional weight updates in PEFT that mirroring the human brain’s ability to store specific skills in distinct, specialized regions.
[…]
  • Introduces sample-based multiplication to realize the efficient SBoRA-FA projection-down and SBoRA-FB projection-up transformations that reduce computational complexity by eliminating one matrix multiplication operation.
2 Related Work
In deep learning, fine-tuning is a fundamental concept that enables leveraging pre-trained models as a starting point for new tasks [7, 9]. This approach is based on the idea that a model trained on one task can be adapted to perform well on another related task.
2.1 Full Fine-Tuning (FFT)
FFT involves updating all of the parameters of a pre-trained model on a new dataset for a specific task as shown in Fig. 1(a). This process adjusts the model’s weights to fit the new task, leading to improved performance. However, the FFT approach has several limitations. Notably, it necessitates storing and deploying numerous copies of large models tailored to different tasks, which can be prohibitively storage-intensive. Furthermore, FFT is computationally costly, particularly for LLMs, and can be overly time-consuming.
2.2 Parameter-Efficient Fine-Tuning (PEFT)
To address the limitations of FFT, researchers have proposed PEFT techniques [9]. These methods adapt only a small subset of a model’s parameters to a new task, aiming to achieve comparable performance to FFT while reducing computational and memory requirements. Early approaches include the use of adapter layers, which add small, trainable modules to each Transformer layer. Another approach is prefix tuning [15], which optimizes a few continuous "prefix" vectors prepended to the input sequence. Both of these approaches significantly reduce the memory and computation needed to fine-tune large models. However, they also have their limitations, such as extra inference latency in the case of adapter layers, and difficulty in optimization for prefix tuning.
2.3 Low-Rank Adaptation (LoRA)
LoRA [10], a prominent PEFT method, has garnered significant attention due to its effectiveness and efficiency. As depicted in Fig. 1(b), the fundamental concept underlying LoRA involves approximating full-rank updates with low-rank updates within the FFT domain. Specifically, given a pre-trained parameter matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, LoRA utilizes two low-rank matrices, $\mathbf{A}\in\mathbb{R}^{r\times k}$ and $\mathbf{B}\in\mathbb{R}^{d\times r}$, to compute the weight update $\mathrm{\Delta}\mathbf{W}\in\mathbb{R}^{d\times k}$. The output vector $\mathbf{h}\in\mathbb{R}^{d\times 1}$ of a LoRA linear layer with input $\mathbf{x}\in\mathbb{R}^{k\times 1}$ can be expressed as
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\mathrm{\Delta}\mathbf{Wx}=\mathbf{W}_{0}\mathbf{x}+\mathbf{BAx}$ (1)
LoRA ensures $\mathrm{\Delta}\mathbf{W}$ is initialized to zero by initializing $\mathbf{A}$ with a uniform distribution and $\mathbf{B}$ with zero. The low-rank decomposition of $\mathrm{\Delta}\mathbf{W}$ into $\mathbf{BA}$ indicates that the rank of $\mathrm{\Delta}\mathbf{W}$ is at most $r$, significantly lower than full-rank updating in FFT, with $r$ being much smaller than the minimum of $d$ and $k$.
LoRA’s low-rank updating approach has demonstrated comparable performance to full-rank updating in tasks like text classification and instruction tuning. It reduces memory usage and computational requirements by decreasing the number of trainable parameters needed. LoRA injects low-rank matrices directly into existing layers without adding extra layers, resulting in no additional inference latency. Switching out different LoRA modules is easy, simplifying deployment in multi-task scenarios. Empirical studies show that LoRA can match or exceed FFT performance using only a fraction of the total parameters, reducing the need for high-performance hardware.
2.4 Variants of LoRA
LoRA has sparked significant research in PEFT methods, including its own variants such as QLoRA [4], QA-LoRA [29], LongLoRA [3], S-LoRA [23], LQ-LoRA [6], MultiLoRA [27], LoRA-FA [32], Tied-LoRA [22], GLoRA [2], and DoRA [17]. QLoRA is an industry-standard technique for PEFT of LLMs, it employs 4-bit quantization on pretrained weights and trains LoRA modules on this quantized representation. Techniques like 4-Bit NormalFloat (NF4) Format, Double Quantization, and Paged Optimizers further minimize memory usage. QA-LoRA reduces the computational burden with group-wise quantization. LongLoRA enables fine-tuning for longer context lengths using sparse local attention. S-LoRA presents a scalable strategy for deploying multiple LoRA modules efficiently. LQ-LoRA refines the quantization scheme for improved performance. MultiLoRA handles complex multi-task learning, LoRA-FA reduces memory overhead, and Tied-LoRA leverages weight tying. GLoRA adapts both weights and activations, and DoRA decomposes weights for enhanced learning capacity. Recently, Orthogonal Adaptation enables efficient merging of customized models without sacrificing fidelity or incurring additional computational costs.
3 Standard Basis Low-Rank Adaptation (SBoRA)
SBoRA, inspired by Orthogonal Adaptation [20], utilizes a predefined orthogonal basis $\mathbf{O}_{A}$ to generate orthogonal subspaces. These subspaces, represented as low-rank projection-down matrices $\mathbf{A}_{\mathit{i}}$, enable independent LoRA fine-tuning for multi-concept customization in diffusion models. The linear layer combining $c$ custom concepts, given a pre-trained weight $\mathbf{W}_{0}$, can be expressed as
$Linear(\mathbf{x})=(\mathbf{W}_{0}+\sum_{i=1}^{c}\lambda_{i}\cdot\mathbf{B}_{i}\mathbf{A}_{i})\mathbf{x}$ (2)
where $i$ denotes the index of the $i$-th LoRA, and $\lambda_{i}$ are scalar factors determined through empirical tuning. The input is a column vector $\mathbf{x}\in\mathbb{R}^{k\times 1}$. $\mathbf{A}_{i}\in\mathbb{R}^{r\times k}$ matrices are constructed by selecting $r$ non-overlapping orthogonal basis vectors from the shared orthogonal basis $\mathbf{O}_{\mathit{A}}\in\mathbb{R}^{k\times k}$. $\mathbf{O}_{\mathit{A}}$ consists of $k$ orthogonal basis vector $\mathbf{o}_{p}\in\mathbb{R}^{1\times k}$, where $p=1,2,\ldots,k$. It can be represented as
$\mathbf{O}_{\mathit{A}}=\left[\begin{array}[]{c}\mathbf{o}_{1}\\ \mathbf{o}_{2}\\ \vdots\\ \mathbf{o}_{k}\end{array}\right]\mathrm{\ with\ }\mathbf{o}_{p}\mathbf{o}_{q}^{T}=0\mathrm{\ if\ }p\neq q$ (3)
To maintain orthogonality between projection-down matrices, these basis vectors must be non-overlapping from $\mathbf{O}_{A}$, ensuring $\mathbf{A}_{p}\mathbf{A}_{q}^{T}=0$ if $p\neq q$. And to minimize crosstalk, $\mathbf{B}_{j}\mathbf{A}_{j}\mathbf{x}=\mathbf{0}$ if $j\neq i$ when $\mathbf{x}$ belongs to subspace $\mathbf{A}_{i}$. During fine-tuning, the orthogonal matrices $\mathbf{A}_{i}$ are frozen while updating the projection-up matrices $\mathbf{B}_{i}$ to learn different concepts.
Figure 2: The diagram illustrates the regional weight update process of SBoRA, showcasing distinct $\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W}$ computing procedures of SBoRA-FA(upper) and SBoRA-FB (lower). The diagram employs different colors to represent frozen, trainable, and zero parameters.
3.1 Orthogonal Standard Basis
The orthogonal adaptation approach offers not only effective multi-concept merging but also memory efficiency. During LoRA fine-tuning, the $\mathbf{A}_{i}$ matrices are frozen, which significantly reduces memory requirements for gradient and intermediate activation storage. This is similar to the LoRA-FA [32], where projection-down matrix $\mathbf{A}$ is randomly initialized and then frozen during training. However, we can further improve this orthogonal adaptation idea by utilizing a standard basis with one-hot vectors as basis vectors. In this case, the orthogonal basis $\mathbf{O}_{A}$ becomes a $k\times k$ identity matrix $\mathbf{I}$. Specifically, $\mathbf{O}_{A}$ can be represented as:
$\mathbf{O}_{A}=\mathbf{I}=\left[\begin{array}[]{c}\mathbf{e}_{1}\\ \mathbf{e}_{2}\\ \vdots\\ \mathbf{e}_{k}\end{array}\right]=\left[\begin{array}[]{cccc}1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&1\end{array}\right]$ (4)
where $\mathbf{e}_{p}\mathbf{e}_{q}^{T}=0$ if $p\neq q$. The standard basis vector $\mathbf{e}_{k}\in\mathbb{R}^{1\times k}$ are one-hot row vectors, each having a single non-zero entry of 1 at index $p$. For instance, $\mathbf{e}_{p}$ can be represented as
[…]
$\uparrow$
$\ p$
where the non-zero component with value one at the index of $p$. In SBoRA-FA, we don’t need to randomly pre-generate an orthogonal-basis matrix $\mathbf{O}_{A}$. Instead, we randomly select $r$ non-overlapping indices between 1 and $k$ to construct a standard subspace basis matrix $\mathbf{A}_{sb}$ with $r$ standard basis row vectors of $\mathbf{e}_{p}$. This approach eliminates the need to store weights for the projection-down matrix, as only the indices of the standard basis vectors are required to represent $\mathbf{A}_{sb}$. The basic structure of SBoRA-FA is depicted in Fig. 1(c), where both $\mathbf{W}_{0}$ and $\mathbf{A}_{sb}$ are frozen, and only the matrix $\mathbf{B}$ is updated with initial values of zero that ensure that the pre-trained models do not alter the model prediction before fine-tuning. Similarity, SBoRA-FB freezes the projection-up as standard-basis low-rank matrix $\mathbf{B}_{sb}$ and only update the projection-down matrix $\mathbf{A}$ during the fine-tuning, in which orthogonal basis matrix $\mathbf{O}_{B}$ is a $d\times d$ identity matrix $\mathbf{I}$, the standard basis vectors are one-hot column vectors instead of row vectors.
$\mathbf{O}_{B}=\mathbf{I}=\left[\begin{array}[]{cccc}\mathbf{e}_{1}&\mathbf{e}_{2}&\cdots&\mathbf{e}_{d}\end{array}\right]=\left[\begin{array}[]{cccc}1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&1\end{array}\right]$ (6)
where $\mathbf{e}_{p}^{T}\mathbf{e}_{q}=0$ if $p\neq q$. The standard basis vector $\mathbf{e}_{p}\in\mathbb{R}^{d\times 1}$ are one-hot column vectors, each having a single non-zero entry of 1 at index $p$. For instance, $\mathbf{e}_{p}$ can be represented as
[…]
$\uparrow$
$\ p$
The basic structure of SBoRA-FB is depicted in Fig. 1(d), where both $\mathbf{W}_{0}$ and $\mathbf{B}_{sb}$ are frozen, and only the matrix $\mathbf{A}$ is updated with initial values of zero that ensure that the pre-trained models do not alter the model prediction before fine-tuning.
3.2 Regional Weight Update
An interesting property of the proposed SBoRA is that the merged weight matrix $\mathbf{W}^{\prime}$ of the fine-tuned model is only regionally updated, with most of the weights remaining unchanged from the original pre-trained weight $\mathbf{W}_{0}$ as depicted in Fig. 2. Specifically, the fine-tuned weight $\mathbf{W}^{\prime}$ of SBoRA-FA can be represented as:
$\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W}=\mathbf{W}_{0}+\mathbf{BA}_{sb}$ (8)
In which, the update matrix $\mathrm{\Delta}\mathbf{W}=\mathbf{BA}_{sb}$ is very sparse, with most of the columns having zero weights due to the one-hot nature of the standard basis subspace matrix $\mathbf{A}_{sb}$ as shown in the upper part of Fig. 2. For example, when $r=2$ and $k=4$ with $\mathbf{A}_{sb}=[\mathbf{e}_{1}\ \mathbf{e}_{4}]^{T}$, the $\mathrm{\Delta}\mathbf{W}$ will only have two non-zero column as
$\mathrm{\Delta}\mathbf{W}=\mathbf{BA}_{sb}=\left[\begin{array}[]{cc}b_{11}&b_{12}\\ b_{21}&b_{22}\\ b_{31}&b_{32}\\ b_{41}&b_{42}\end{array}\right]\left[\begin{array}[]{cccc}1&0&0&0\\ 0&0&0&1\end{array}\right]=\left[\begin{array}[]{cccc}b_{11}&0&0&b_{12}\\ b_{21}&0&0&b_{22}\\ b_{31}&0&0&b_{32}\\ b_{41}&0&0&b_{42}\end{array}\right]$ (9)
The fine-tuned weight $\mathbf{W}^{\prime}$ is given by
$\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W}=\left[\begin{array}[]{cccc}w_{11}+b_{11}&w_{12}&w_{13}&w_{14}+b_{12}\\ w_{21}+b_{21}&w_{22}&w_{23}&w_{24}+b_{22}\\ w_{31}+b_{31}&w_{32}&w_{33}&w_{34}+b_{32}\\ w_{41}+b_{41}&w_{42}&w_{43}&w_{44}+b_{42}\end{array}\right]$ (10)
For SBoRA-FB, when $r=2$ and $d=4$ with $\mathbf{B}_{sb}=[\mathbf{e}_{1}\ \mathbf{e}_{4}]$, the $\mathrm{\Delta}\mathbf{W}$ will only have two non-zero row as
$\mathrm{\Delta}\mathbf{W}=\mathbf{B}_{sb}\mathbf{A}=\left[\begin{array}[]{cc}1&0\\ 0&0\\ 0&0\\ 0&1\end{array}\right]\left[\begin{array}[]{cccc}b_{11}&b_{12}&b_{13}&b_{14}\\ b_{12}&b_{22}&b_{23}&b_{24}\end{array}\right]=\left[\begin{array}[]{cccc}b_{11}&b_{12}&b_{13}&b_{14}\\ 0&0&0&0\\ 0&0&0&0\\ b_{21}&b_{22}&b_{23}&b_{24}\end{array}\right]$ (11)
The fine-tuned weight $\mathbf{W}^{\prime}$ is given by
$\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W}=\left[\begin{array}[]{cccc}w_{11}+b_{11}&w_{12}+b_{12}&w_{13}+b_{13}&w_{14}+b_{14}\\ w_{21}&w_{22}&w_{23}&w_{24}\\ w_{31}&w_{32}&w_{33}&w_{34}\\ w_{41}+b_{21}&w_{42}+b_{22}&w_{43}+b_{23}&w_{44}+b_{24}\end{array}\right]$ (12)
In these examples, only half of the weight matrix $\mathbf{W}^{\prime}$ updated, specifically two columns or two rows $(r=2)$, while the remaining weights remain unchanged. In practice, $r\ll$ min$(k,d)$, therefore only a small portion of the original weights will be updated. More details of the regional weight update property of SBoRA-FA and SBoRA-FB can be visualized in Fig. 2. It resembles the localized learning process observed in neuroscience, where specific cognitive functions are associated with distinct brain regions. SBoRA efficiently adapts to new tasks by updating specific rows of the weight matrix, preserving existing knowledge, similar to how the brain reorganizes and refines neural connections in response to new experiences.
4 Complexity Analysis of SBoRA
The implementation of SBoRA offers a high degree of flexibility, and we propose sampling-based multiplication, sampleMul, to achieve lower memory and computational requirements. Essentially, both SBoRA-FA/FB can be implemented similarly to conventional LoRA through matrix multiplication, with just frozen the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$. As a result, the computational complexity and memory usage of SBoRAs are similar to those of LoRA. However, we can optimize memory usage by representing the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$ as a $r$-dimensional vector with integer coefficients corresponding to the indices of the selected subspace basis vectors. The basis vector indices are initialized by randomly sampling $r$ unique numbers between 1 and $k$ for SBoRA-FA or between 1 and $d$ for SBoRA-FB. This implementation of SBoRA achieves a significant reduction in parameter storage, approximately 50%, while also saving around half of the gradient storage due to the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$ being frozen. Table 1 provides an analysis of the memory requirements for SBoRA compared to LoRA.
Table 1: Analysis of extra memory storage bring by a single LoRA module and SBoRA-FA/FB. Through the utilization of a standard basis and index-based updates SBoRA enables the replacement of an entire matrix with a 1-dimensional vector. This reduction leads to a significant decrease of approximately 50% total parameters.
Method Trainable Params Total Params Gradient Memory
LoRA $(k+d)\times r$ $(k+d)\times r$ $(k+d)\times r$
SBoRA-FA $d\times r$ $r+k\times r$ $d\times r$
SBoRA-FB $k\times r$ $r+d\times r$ $k\times r$
To reduce computational complexity, we can leverage one-hot standard basis vectors to simplify the projection transformation, effectively converting it into a component sampling process that eliminates the need for multiplication operations. Specifically, we select components of the input vector based on the indices of the standard basis vectors. This allows us to sample the input vector into the projection-down vector based on the corresponding indices in $\mathbf{A}_{sb}$. The forward process of SBoRA-FA with sampleMul can be expressed as:
[…]
where $\mathbf{x}[seq,\mathbf{A}_{sb}]$ is a component sampling process method for implementing $\mathbf{A}_{sb}\mathbf{x}$. This approach is also applied to the case of SBoRA-FB in the projection-up transformation. We can utilize the $index\_add$($\cdot$) operation to incorporate the non-zero values from the output projection-up vector into the pre-trained weight output vector ($\mathbf{W}_{0}\mathbf{x}$) based on the corresponding indices in $\mathbf{B}_{sb}$. The forward process of SBoRA-FB with sampleMul can be expressed as:
[…]
Notably, for SBoRA-FB, projection-down vector $\mathbf{z}$ is directly added into the pre-trained weight output vector through index addition method, eliminating the need for matrix multiplication operations. During this process, two matrix multiplication operations $\mathbf{B}(\mathbf{A}_{sb}\mathbf{x})$ / $\mathbf{Ax}$ and $\mathbf{W}_{0}\mathbf{x}$, and one matrix addition operation of $\mathbf{W}_{0}\mathbf{x}+\mathbf{BA}_{sb}\mathbf{x}$ / $\mathbf{W}_{0}\mathbf{x}+\mathbf{B}_{sb}\mathbf{Ax}$ are involved.
It reduces computation complexity by saving one matrix multiplication operation compared to LoRA[10]. For our experiments, we used the SBoRA sampleMul implementation due to its lower computational complexity and memory usage compared to traditional LoRA methods. Table 2 provides an analysis of computational requirements for SBoRA using sampleMul compared to LoRA, assuming input $\mathbf{x}$ in the shape of $[d\times 1]$.
Table 2: Computational analysis and comparison of LoRA and SBoRA’s sampleMul implementation, with a pre-trained weight shape of $k\times d$ and a LoRA rank of $r$. Notably, in the SBoRA_sampleMul implementations, the number of multiplication and addition operations is significantly reduced.
Method Multiplication Addition
LoRA  $k\times d$; $(k+d)\times r$  $k\times(d-1)$; $r\times(d-1)+k\times(r-1)$; $1\times k$
SBoRA-FA $k\times d$; $k\times r$ $k\times(d-1)$; $k\times(r-1)$; $1\times k$
SBoRA-FB $k\times d$; $d\times r$ $k\times(d-1)$; $r\times(d-1)$; $1\times r$
5 Experiment
We conducted a comprehensive evaluation of SBoRA’s effectiveness across a range of tasks. To begin with, we compared the performance of SBoRA-FA and SBoRA-FB to that of several other well-known PEFT methods by fine-tuning LLMs, specifically LLaMA-7B, on both commonsense reasoning tasks and arithmetic reasoning tasks. Additionally, given the recent release of LLaMA3 [18], we expanded our evaluation to assess SBoRA’s performance using LLaMA3-8B. Furthermore, recognizing the growing interest in quantized models, we also evaluated SBoRA on the quantized versions of LLaMA-7B LLaMA-13B and LLaMA3-8B models, utilizing the techniques introduced in QLoRA [4].
5.1 Evaluating SBoRA on Commonsense Reasoning Tasks
We evaluated SBoRA-FA/FB against LoRA, and DoRA approach on the LLaMA-7B and LLaMA3-8B language models for commonsense reasoning tasks. We followed the evaluation method introduced in DoRA [17] and LLM-Adapter [11] which used a comprehensive training dataset created by aggregating eight tasks, evaluating models on individual testing datasets for each task. In addition, the results of GPT-3.5 are included for reference as shown in Table 3.
For a comprehensive comparison, we initially fine-tuned models with SBoRA-FA and SBoRA-FB, both using a double rank of 64 to maintain a nearly identical number of trainable parameters (TP) as LoRA (rank=32), with only additional storage for standard basis indices. We then conducted experiments with rank 32 for SBoRA-FA/FB and 64 for other baselines to investigate if SBoRA could achieve better performance with half the trainable parameters. Consistency was maintained by using one training epoch and learning rate of 2e-4 for all experiments.
In LLaMA-7B, SBoRA-FA (rank=64) outperforms all baseline methods with a similar number of trainable parameters, including LoRA (rank=32), and DoRA (rank=32). SBoRA-FA (rank=32) achieves commendable performance, matching the accuracy of DoRA (rank=32) and trailing behind the best accuracy of LoRA (rank=32) by only 1.1%. Despite having approximately half the trainable parameters, SBoRA performs comparably to models with the same rank. While in LLaMA3-8B, the top two results are achieved by SBoRA-FA with rank 64 and 32, respectively.
Importantly, under identical settings and rank with NVIDIA Geforce RTX 4090, SBoRA-FA/FB significantly reduces training time by half compared to DoRA. Fine-tuning with DoRA takes around 7-8 hours, while achieving comparable or better performance with SBoRA can be done in just 3-4 hours. The notable computational efficiency of SBoRA saves a substantial amount of time.
Table 3: Comparison of LLaMA-7B and LLaMA3-8B with different PEFT methods, and evaluating on the commonsense reasoning task. The first row list the results of GPT-3.5 for reference. We report the accuracy (%) results for each of the eight sub-tasks as well as the average accuracy (%), higher is better for all metrics. The column headers indicates TP for the number of trainable parameters and $r$ for the rank.
Model Method $r$ TP BoolQ PIQA SIQA HS WG ARC-e ARC-c OBQA Avg
GPT-3.5 - - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA-7B LoRA 32 56.1M 66.8 81.1 78.4 53.5 80.5 81.1 61.9 79.4 72.8
DoRA 57.0M 68.8 82.0 70.6 57.6 73.2 79.4 64.2 78.2 71.7
SBoRA-FA 28.0M 68.0 79.7 76.2 54.4 79.1 79.8 61.3 75.0 71.7
SBoRA-FB 28.0M 66.1 64.2 74.8 57.2 71.5 80.4 62.4 75.8 64.9
LoRA 64 112.2M 62.1 81.8 78.2 62.9 78.6 79.8 63.7 81.2 73.5
DoRA 113.1M 68.7 82.8 78.2 64.8 62.9 79.7 64.8 80.0 72.7
SBoRA-FA 56.1M 68.2 81.3 77.6 74.7 81.1 80.8 62.8 79.4 75.7
SBoRA-FB 56.1M 66.5 79.2 76.7 59.2 76.5 76.8 59.0 74.4 71.0
LLaMA3-8B LoRA 32 56.6M 71.9 86.7 80.4 94.0 85.6 87.8 75.9 83.6 83.2
DoRA 57.4M 73.6 87.1 80.8 94.4 86.1 88.8 78.3 84.2 84.2
SBoRA-FA 25.2M 73.3 87.8 79.1 93.9 85.2 89.9 80.0 86.0 84.4
SBoRA-FB 31.5M 72.9 86.3 78.8 92.6 83.0 88.8 76.3 85.0 83.0
LoRA 64 113.2M 72.5 87.8 80.3 94.4 86.4 88.7 79.3 85.2 84.3
DoRA 114.0M 70.5 86.0 80.3 91.8 83.7 86.2 74.7 83.2 82.1
SBoRA-FA 50.3M 74.0 88.3 80.8 94.3 86.3 89.9 78.7 86.6 84.9
SBoRA-FB 62.9M 71.8 85.2 79.2 91.4 82.9 86.7 74.0 83.4 81.8
5.2 Evaluating SBoRA on Arithmetic Reasoning
We also evaluated the effectiveness of SBoRA on the Arithmetic Reasoning task using the LLaMA-7B and LLaMA3-8B models. The fine-tuning and evaluation datasets are constructed following the dataset settings in LLM-Adapters [11]. We performed tests with ranks 32 and 64. As the training set for arithmetic reasoning is smaller (10k) compared to the commonsense reasoning training set, we increased the number of training epochs to 3 for better convergence and set the learning rate to 3e-4.
Table 4 presents the performance of PEFT methods, including GPT-3.5. On average, for LLaMA-7B, GPT-3.5 (175B) outperforms adapter-based PEFT LLMs in accuracy. However, for simpler math reasoning tasks like MultiArith, adapter-based methods outperform GPT-3.5, and SBoRA-FA achieves the best performance. Initially, both SBoRA-FA and SBoRA-FB used a rank of 64 for consistency in trainable parameters. SBoRA-FA achieves the best average accuracy. Increasing the rank of other baselines to 64 improves accuracy but still lags behind SBoRA-FA. Reducing SBoRA’s rank to 32 maintains comparable performance to rank-32 baselines. Notably, for the challenging task AQuA, SBoRA-FA with rank 32 excels. Moving to LLaMA3-8B, SBoRA-FA with rank 32 achieves the best performance, followed by SBoRA-FA with rank 64.
SBoRA-FA/FB achieves comparable or superior results to other baseline PEFT methods with half the trainable parameters for the same rank. Arithmetic reasoning tasks are particularly well-suited for SBoRA due to the specialized knowledge required in mathematics. SBoRA’s regional weight update property makes it more suitable for higher ranks like 64, impacting more pre-trained weight regions.
Table 4: Comparison of LLaMA-7B and LLaMA3-8B with different PEFT methods, and evaluating on the arithmetic reasoning task. The first row list out the results of GPT-3.5 for reference. We report the accuracy results for each of the sub-tasks as well as the average accuracy, higher is better for all metrics. The number of trainable parameters (TP) can be found in each row.
LLM Method Rank TP MultiArith GSM8K AddSub AQuA SingleEq SVAMP Avg
GPT-3.5 - - - 83.8 56.4 85.3 38.9 88.1 69.9 70.4
LLaMA-7B LoRA 32 56.1M 94.5 36.3 81.8 15.0 82.7 45.6 59.3
DoRA 57.0M 95.7 36.2 78.7 15.4 81.7 46.6 59.1
SBoRA-FA 28.0M 95.5 34.6 79.7 20.1 78.9 44.8 58.9
SBoRA-FB 28.0M 92.2 31.0 77.5 15.7 78.5 41.8 56.1
LoRA 64 112.2M 94.0 36.8 84.3 17.3 82.3 44.7 59.9
DoRA 113.1M 95.0 35.5 84.1 20.1 85.0 47.1 61.1
SBoRA-FA 56.1M 97.8 36.6 85.1 19.3 83.9 48.5 61.9
SBoRA-FB 56.1M 94.8 33.1 77.5 16.9 78.5 40.6 56.9
LLaMA3-8B LoRA 32 56.6M 68.3 50.5 83.3 35.8 87.2 71.2 66.1
DoRA 57.4M 97.3 62.0 90.9 25.6 94.9 73.4 74.0
SBoRA-FA 25.2M 99.5 66.0 91.9 30.3 97.4 75.8 76.8
SBoRA-FB 31.5M 98.0 57.2 92.2 33.9 94.1 69.6 74.2
LoRA 64 113.2M 97.2 56.3 92.7 22.8 92.3 69.3 71.8
DoRA 114.0M 97.8 55.2 91.1 24.0 94.7 72.0 72.5
SBoRA-FA 50.3M 99.2 64.7 94.4 24.8 98.0 75.0 76.0
SBoRA-FB 62.9M 98.2 50.9 87.1 28.0 91.7 63.0 69.8
5.3 QSBoRA Evaluation on MMLU
We explored the effectiveness of SBoRAs in the memory-efficient QLoRA framework [4]. QLoRA utilizes techniques including 4-bit NormalFloat (NF4) double quantization, gradient checkpointing, and paged optimizer to reduce memory usage. Following the experiments in QLoRA, we fine-tuned the NF4 quantized versions of LLaMA-7B/13B and LLaMA3-8B using the Alpaca [24] and sampled FLAN v2 [28] datasets. Model performance was evaluated on the MMLU benchmark [8] by comparing the mean 5-shot MMLU test accuracy.
To ensure fairness, we used the same hyperparameter settings for all experiments. For the Alpaca dataset, we applied a batch size (bs) of 16 and 10k training steps. To maintain a similar number of training epochs with Alpaca, for FLAN v2, we used batch size 8 and 20k training steps. In the QLoRA framework, we compared SBoRA with LoRA and DoRA, using rank 64 for all methods. Adapters were added to linear layers in both self-attention and MLP modules.
The results shown in Table 5 reveal that within LLaMA-7B, both QSBoRA-FA/FB achieve comparable accuracy results to QLoRA. In the case of LLaMA-13B and LLaMA3-8B, both QSBoRA-FA/FB outperform QLoRA and QDoRA in terms of accuracy, indicating its compatibility with larger-scale models.
Table 5: Finetune LLaMA-7B/13B and LLaMA3-8B on Alpaca and Flan v2. We measured the performance using MMLU benchmark and report the 5-shot average accuracy. The training settings and the number of trainable parameters (TP) are included.
NFloat4 LLaMA-7B LLaMA-13B LLaMA3-8B
PEFT TP Alpaca Flanv2 TP Alpaca Flanv2 TP Alpaca Flanv2
QLoRA 80.0M 37.9 44.4 125.2M 45.4 46.7 83.9M 51.9 49.5
QDoRA 80.6M 38.0 42.8 126.2M 46.7 48.8 84.6M 53.0 51.9
QSBoRA-FA 43.5M 36.5 43.1 68.2M 49.0 51.0 44.0M 56.5 56.4
QSBoRA-FB 36.4M 36.9 43.4 57.0M 48.3 50.5 39.8M 54.5 55.0
6 Conclusion
We introduce SBoRA, a novel approach for parameter-efficient fine-tuning of LLMs. Our proposed SBoRA-FA and SBoRA-FB methods utilize the low-rank projection-down and projection-up matrices constructed from standard basis vectors, leveraging one-dimensional index representations and sample-based multiplication. This innovative approach reduces memory and computational requirements while enhancing learning capacity, outperforming LoRA and matching the performance of DoRA. Empirical results demonstrate that SBoRA-FA is the preferred method for single-task LLM fine-tuning, conserving activation memory while achieving superior performance.
The SBoRA approach enables regional weight updates, preserving most of the pre-trained model weights while efficiently adapting to new tasks. This localized learning process draws parallels with the modular organization of the brain, where distinct cognitive functions are localized to specific brain regions. This analogy highlights the potential of SBoRA to inspire AI architectures that mimic the efficiency and adaptability of biological neural systems.
SBoRA holds immense potential for further development, particularly in multi-task training. The introduction of Multi-SBoRA would create a powerful framework for efficient adaptation to multiple tasks. Each task would undergo independent, non-overlapping weight fine-tuning with SBoRA, allowing the integration of task-specific knowledge while minimizing interference and maximizing the model’s capacity to leverage shared information. This approach enables the maintenance of distinct capabilities for each task within a single model, paving the way for more efficient and effective AI systems